# An Adaptive Multi-Agent Bitcoin Trading System

### v1

@@ 4 Evaluation Methodology
The system was evaluated on daily Bitcoin data from July 2024 to April 2025. Market regimes were defined using moving average trends to distinguish bullish, bearish, and sideways phases. Performance was measured against baselines including buy-and-hold BTC, a Nasdaq Crypto Index portfolio, and an equal-weighted basket of the top 30 cryptocurrencies. Key evaluation metrics included cumulative returns, Sharpe ratio, prediction accuracy, and regret.
@@ 5 Results and Discussion
Regime Metric Quants (w/o LTF) Decision (w/o LTF) Signals (w/o LTF) Quants (LTF) Decision (LTF) Signals (LTF) Baseline All Periods Total Return (%) 21.05 11.50 21.14 6.73 14.51 23.92 18.27 Daily Return (mean $\pm$ std) 0.08 $\pm$ 1.582 0.052 $\pm$ 1.630 0.08 $\pm$ 1.609 0.035 $\pm$ 1.519 0.059 $\pm$ 1.524 0.088 $\pm$ 1.592 0.069 $\pm$ 1.379 Sharpe Ratio 0.0504 0.0318 0.0499 0.0227 0.0388 0.0553 0.0497 Accuracy 0.6297 0.6357 0.6358 0.6165 0.6402 0.6467 Sideways 1/2 Total Return (%) -2.02 -3.38 2.99 0.08 1.35 7.13 -0.30 Daily Return (mean $\pm$ std) -0.002 $\pm$ 1.417 -0.053 $\pm$ 1.602 0.005 $\pm$ 1.521 0.01 $\pm$ 1.439 -0.005 $\pm$ 1.429 0.031 $\pm$ 1.436 -0.011 $\pm$ 1.366 Sharpe Ratio -0.0014 -0.0333 0.0033 0.0068 -0.0033 0.0216 -0.0081 Accuracy 0.6533 0.6472 0.6377 0.6561 0.6809 0.6910 Bullish Total Return (%) 36.05 33.34 30.92 21.48 25.48 29.67 27.27 Daily Return (mean $\pm$ std) 0.458 $\pm$ 1.655 0.456 $\pm$ 1.587 0.415 $\pm$ 1.610 0.33 $\pm$ 1.463 0.374 $\pm$ 1.521 0.396 $\pm$ 1.637 0.378 $\pm$ 1.270 Sharpe Ratio 0.2765 0.2873 0.2578 0.2252 0.2457 0.2417 0.2972 Accuracy 0.6249 0.6313 0.6427 0.5921 0.6177 0.6039 Sideways 2 Total Return (%) 6.21 3.01 5.63 9.08 5.83 3.73 5.91 Bearish Total Return (%) -18.82 -20.98 -17.66 -23.43 -17.66 -15.87 -14.00 Daily Return (mean $\pm$ std) -0.151 $\pm$ 1.701 -0.18 $\pm$ 1.652 -0.128 $\pm$ 1.701 -0.207 $\pm$ 1.648 -0.154 $\pm$ 1.618 -0.134 $\pm$ 1.719 -0.116 $\pm$ 1.445 Sharpe Ratio -0.0889 -0.1088 -0.0750 -0.1254 -0.0949 -0.0778 -0.0805 Accuracy 0.6033 0.6266 0.6298 0.5852 0.6020 0.6163
Empirical evaluation on Bitcoin data (July 2024–April 2025) shows that the system consistently improved on static baselines across market regimes. Without long-term (weekly) feedback, the quantitative agent, as compared to the static buy and hold BTC portfolio, delivered over 30% higher returns in bullish phases and a 15% overall improvement, while the signals agent turned sideways markets from a slight loss into a gain of more than 100%. Introducing long-term weekly feedback amplified these effects: the signals agent achieved a 31% overall improvement as compared to without LTF, boosted sideways performance, and reduced bearish losses by more than 10%. These gains demonstrate that verbal feedback can serve as a scalable alternative to fine-tuning, enhancing both profitability and robustness under regime shifts.
@@
Crucially, the verbal feedback mechanism was found to be decisive. Daily reflections reduced repeated short-term errors, while weekly feedback stabilized long-term strategy formation. Together, they allowed the agents to adapt their focus dynamically, replicating the benefits of human oversight without additional training costs.
@@ 6.1 Daily Feedback without Weekly (Long Term) Feedback
On 4 November 2024, the Quants agent leaned heavily on bearish technicals (MACD -76.84, 70% of prices below VWAP, ADX = 27.33) and recommended a conservative 35% BTC / 65% cash allocation. The Signals agent, however, highlighted bullish catalysts: positive sentiment mean (+0.1164) and strong news flow on miner AI integration and \$200K price targets, leading it to recommend 70% BTC / 30% cash. The Decision agent synthesized these by overweighting Quants’ technical caution, cutting BTC exposure to 40%, which proved overly conservative since Bitcoin actually gained +2.24% that day.
Reflect feedback was explicit: Quants had overemphasized bearish confluence without giving enough weight to neutral RSI/Bollinger signals, while the Decision agent was criticized for failing to quantify the strength of news catalysts (e.g., \$200K targets as asymmetric upside). Signals, by contrast, received positive feedback for identifying the correct bullish call, with only minor suggestions to better calibrate FGI thresholds.
This feedback directly influenced the allocations on 5 November 2024. Quants softened its stance: though MACD remained bearish and ADX weak, it noted VWAP positioning and elevated on-chain activity (\$50.58B volume, 528k addresses) as possible accumulation signals, recommending a higher 60% BTC exposure than the day before. Signals doubled down on its bullish framing, prioritizing the Trump election-driven rally to \$75k and interpreting FGI=70 as a bullish but not extreme greed signal, leading to 65% BTC. The Decision agent, informed by Reflect’s critique of over-hedging, gave greater weight (65%) to Signals’ event-driven catalysts while still acknowledging Quants’ caution, settling on 55% BTC / 45% cash. This trade aligned with the sharp November rally, showing how the feedback loop enabled adaptation: Quants adjusted away from excessive pessimism, Signals refined its interpretation of sentiment extremity, and the Decision agent corrected its tendency to over-hedge, thereby capturing the upside more effectively.
6.2 Daily and Weekly Feedback
@@
The weekly long-term feedback mechanism generates verbal, template-driven guidance at the end of each completed seven-day block. It evaluates performance relative to a passive baseline using regret, Sharpe ratio, and return differentials, and then produces standardised text such as “prioritize high-confidence inputs” or “reconstruct your indicator selection.” Importantly, this feedback does not alter any quantitative parameters: indicator thresholds, weightings, and allocation rules remain fixed. Instead, the generated text is re-introduced into the subsequent week’s prompts as additional context.
This design created an asymmetric effect across agents. The Signals agent, whose reasoning depends on interpreting natural language (news flow, sentiment reports, and textual priors), was able to incorporate such guidance effectively. For example, phrases such as “your signals consistently beat the baseline” reinforced its conviction, while instructions to prioritise high-confidence sources translated directly into more decisive sentiment weighting. This alignment between the form of feedback and the nature of the task explains why Signals exhibited marked improvement, as observed in the weekly narratives where it explicitly referenced feedback when framing its allocation logic.
@@ 7 Contributions and Conclusion
In sum, this thesis shows that an adaptive, feedback-driven multi-agent system can generate consistent Bitcoin trading profits under volatile conditions. By prioritizing profitability over predictive accuracy and employing language-based feedback as its optimization mechanism, the system demonstrates a scalable pathway for deploying LLMs in financial trading.

### v2

@@ 4 Evaluation Methodology
The system was evaluated on daily Bitcoin data from July 2024 to April 2025. Market regimes were defined using moving-average trends to classify days as bullish, bearish, or sideways. Performance was assessed using cumulative returns, Sharpe ratio, prediction accuracy, regret, and aggregated period-level returns.
Prediction accuracy. Accuracy was computed using a soft scoring function that compares the agent’s confidence in the correct class to its highest predicted class. Let $s_{\text{real}}$ denote the predicted probability of the true regime and $s_{\max}$ the maximum predicted probability across regimes. The soft accuracy score is
$\text{Acc}=\begin{cases}0,&s_{\max}=0,\\[4.0pt] \frac{s_{\text{real}}}{s_{\max}},&\text{otherwise}.\end{cases}$
This yields a value of $1$ when the correct regime is the top prediction and a fractional score when the correct regime is predicted with lower confidence.
Daily returns. Daily returns were computed from the agent’s BTC allocation. If $r_{t}^{\text{BTC}}$ is the Bitcoin return for day $t$ and $w_{\text{BTC}}$ is the agent’s BTC weight (expressed as a percentage), the portfolio’s daily return is
$r_{t}=r_{t}^{\text{BTC}}\cdot\frac{w_{\text{BTC}}}{100}.$
Cumulative return. Cumulative return measures the total percentage gain relative to the initial portfolio value. Let $V_{t}$ be the portfolio value at time $t$ and $V_{0}=100$ the initial value. The cumulative return is
$\text{CR}_{t}=\frac{V_{t}-V_{0}}{V_{0}}\times 100.$
Regret. Regret quantifies the performance difference between each agent and the baseline buy-and-hold portfolio. If $V_{t}^{\text{agent}}$ and $V_{t}^{\text{base}}$ denote the agent and baseline portfolio values, regret is defined as
$\text{Regret}_{t}=V_{t}^{\text{agent}}-V_{t}^{\text{base}}.$
Sharpe ratio. Sharpe ratio was computed using the empirical mean and standard deviation of daily returns for each agent. Let $\mu$ and $\sigma$ denote the sample mean and standard deviation of returns. The Sharpe ratio is
$\text{Sharpe}=\begin{cases}0,&\sigma=0,\\[4.0pt] \frac{\mu}{\sigma},&\text{otherwise}.\end{cases}$
Weekly and monthly aggregated returns. For longer horizons, returns were averaged across all days in the corresponding week or month. Given a set of returns $\{r_{1},r_{2},\ldots,r_{n}\}$, the aggregated return is
$\overline{r}=\frac{1}{n}\sum_{i=1}^{n}r_{i}.$
@@ 5 Results and Discussion
Regime Metric Quants (w/o LTF) Decision (w/o LTF) Signals (w/o LTF) Quants (LTF) Decision (LTF) Signals (LTF) Baseline All Periods Total Return (%) 21.05 11.50 21.14 6.73 14.51 23.92 18.27 Daily Return (mean $\pm$ std) 0.08 $\pm$ 1.582 0.052 $\pm$ 1.630 0.08 $\pm$ 1.609 0.035 $\pm$ 1.519 0.059 $\pm$ 1.524 0.088 $\pm$ 1.592 0.069 $\pm$ 1.379 Sharpe Ratio 0.0504 0.0318 0.0499 0.0227 0.0388 0.0553 0.0497 Accuracy 0.6297 0.6357 0.6358 0.6165 0.6402 0.6467 Sideways Total Return (%) -2.02 -3.38 2.99 0.08 1.35 7.13 -0.30 Daily Return (mean $\pm$ std) -0.002 $\pm$ 1.417 -0.053 $\pm$ 1.602 0.005 $\pm$ 1.521 0.01 $\pm$ 1.439 -0.005 $\pm$ 1.429 0.031 $\pm$ 1.436 -0.011 $\pm$ 1.366 Sharpe Ratio -0.0014 -0.0333 0.0033 0.0068 -0.0033 0.0216 -0.0081 Accuracy 0.6533 0.6472 0.6377 0.6561 0.6809 0.6910 Bullish Total Return (%) 36.05 33.34 30.92 21.48 25.48 29.67 27.27 Daily Return (mean $\pm$ std) 0.458 $\pm$ 1.655 0.456 $\pm$ 1.587 0.415 $\pm$ 1.610 0.33 $\pm$ 1.463 0.374 $\pm$ 1.521 0.396 $\pm$ 1.637 0.378 $\pm$ 1.270 Sharpe Ratio 0.2765 0.2873 0.2578 0.2252 0.2457 0.2417 0.2972 Accuracy 0.6249 0.6313 0.6427 0.5921 0.6177 0.6039 Bearish Total Return (%) -18.82 -20.98 -17.66 -23.43 -17.66 -15.87 -14.00 Daily Return (mean $\pm$ std) -0.151 $\pm$ 1.701 -0.18 $\pm$ 1.652 -0.128 $\pm$ 1.701 -0.207 $\pm$ 1.648 -0.154 $\pm$ 1.618 -0.134 $\pm$ 1.719 -0.116 $\pm$ 1.445 Sharpe Ratio -0.0889 -0.1088 -0.0750 -0.1254 -0.0949 -0.0778 -0.0805 Accuracy 0.6033 0.6266 0.6298 0.5852 0.6020 0.6163
Empirical evaluation on Bitcoin data from July 2024 to April 2025 (Table 2) shows that the system consistently outperformed static baselines across all market regimes.
Effect of Daily Feedback. In the absence of long-term (weekly) feedback, the quantitative agent outperformed the static buy-and-hold Bitcoin portfolio by more than 30% in bullish phases and achieved a 15% improvement overall. The signals agent demonstrated even stronger relative gains: it converted a marginally negative sideways performance into a gain exceeding 100%. The decision agent exhibited the highest Sharpe ratio during bullish periods, but its allocations tended to cluster around a near 50–50 split. This likely reflects difficulties in reconciling potentially conflicting information from the quantitative and signals agents.
Effect of daily and weekly (long-term) feedback. Introducing weekly long-term feedback amplified several of these effects. The signals agent achieved a 31% improvement relative to its performance without long-term feedback, with substantial gains in sideways markets and a reduction of bearish losses by more than 10%. In contrast, both the decision and quantitative agents experienced declines in overall performance. The signals agent, however, benefited considerably from the structure of the prompt injections, which strengthened its ability to adjust its strategy over extended horizons.
Regime-level comparison. During bearish phases, the signals agent without long-term feedback achieved both the highest Sharpe ratio and the highest directional accuracy, although the baseline buy-and-hold strategy naturally produced the smallest negative return. In bullish phases, the quantitative agent without long-term feedback delivered the highest recorded returns. Under sideways conditions, the signals agent with long-term feedback outperformed all others, generating a gain of 7% alongside the highest Sharpe ratio and accuracy.
@@
Interpretation. The verbal feedback mechanism proved to be central to these improvements. Daily reflections mitigated repeated short-term errors, while weekly feedback contributed to more stable long-horizon strategy formation. In combination, these mechanisms enabled the agents to adjust their analytical focus dynamically, delivering many of the benefits associated with human oversight without incurring additional training costs.
Overall, the results indicate that the quantitative agent performs best in bullish regimes, the baseline and signals agent outperform in bearish regimes, and the signals agent dominates in sideways markets. These patterns are examined in greater detail in the subsequent case studies.
@@ 6.1 Daily Feedback without Weekly (Long Term) Feedback
On 4 November 2024 as seen in figure  7, the Quants agent leaned heavily on bearish technicals (MACD -76.84, 70% of prices below VWAP, ADX = 27.33) and recommended a conservative 35% BTC / 65% cash allocation. The Signals agent, however, highlighted bullish catalysts: positive sentiment mean (+0.1164) and strong news flow on miner AI integration and \$200K price targets, leading it to recommend 70% BTC / 30% cash. The Decision agent synthesized these by overweighting Quants’ technical caution, cutting BTC exposure to 40%, which proved overly conservative since Bitcoin actually gained +2.24% that day.
Reflect feedback was explicit: Quants had overemphasized bearish indicators without giving enough weight to neutral RSI/Bollinger signals, while the Decision agent was criticized for failing to quantify the strength of news catalysts (e.g., \$200K targets as asymmetric upside). Signals, by contrast, received positive feedback for identifying the correct bullish call, with only minor suggestions to better calibrate FGI thresholds.
This feedback directly influenced the allocations on 5 November 2024. Quants softened its stance: though MACD remained bearish and ADX weak, it noted VWAP positioning and elevated on-chain activity (\$50.58B volume, 528k addresses) as possible accumulation signals, recommending a higher 60% BTC exposure than the day before. Signals doubled down on its bullish framing, prioritizing the Trump election-driven rally to \$75k and interpreting FGI=70 as a bullish but not extreme greed signal, leading to 65% BTC. The Decision agent, informed by Reflect’s critique of over-hedging, gave greater weight (65%) to Signals’ event-driven catalysts while still acknowledging Quants’ caution, settling on 55% BTC / 45% cash. The agents successfully anticipated the bullish movement: Bitcoin rose by 2.24% that day, and all the three agents generated returns of approximately 5%. This trade aligned with the sharp November rally, showing how the feedback loop enabled adaptation: Quants adjusted away from excessive pessimism, Signals refined its interpretation of sentiment extremity, and the Decision agent corrected its tendency to over-hedge, thereby capturing the upside more effectively.
6.2 Daily and Weekly Feedback
The performance can be observed in figure  7. The weekly long-term feedback mechanism generates verbal, template-driven guidance at the end of each completed seven-day block. It evaluates performance relative to a passive baseline using regret, Sharpe ratio, and return differentials, and then produces standardised text such as “prioritize high-confidence inputs” or “reconstruct your indicator selection.” Importantly, this feedback does not alter any quantitative parameters: indicator thresholds, weightings, and allocation rules remain fixed. Instead, the generated text is re-introduced into the subsequent week’s prompts as additional context.
The following weekly long-term feedback was fixed for all agents throughout the entire week ( 17th nov 2024 - 23rd nov 2024):
QUANTS: Regret shows how often you defeated the returns by baseline, returns shows how much better or worse your results were, and Sharpe ratio shows how much better or worse your risk-adjusted performance was, each calculated as a percentage difference from the baseline. Regret (Outperformance): 100.0. Your strategy dominated the baseline, achieving superior performance almost every day. Continue monitoring indicator efficiency, but avoid unnecessary changes. Sharpe Ratio: -190.96%. Severe volatility. Your allocations are erratic and lead to high losses. Redesign your strategy to prioritize stability. Returns: -178.35%. Severe underperformance. Your market classifications are frequently wrong, leading to heavy losses. Consider redesigning your strategy.
SIGNALS: Regret shows how often you defeated the returns by baseline, returns shows how much better or worse your results were, and Sharpe ratio shows how much better or worse your risk-adjusted performance was, each calculated as a percentage difference from the baseline. Regret (Outperformance): 100.0. Your sentiment signals delivered superior results, consistently beating the baseline. Continue monitoring source quality to ensure sustained performance. Sharpe Ratio: -78.27%. Severe volatility in sentiment. Your signals are misleading and cause erratic performance. Re-evaluate your filtering logic. Returns: -79.26%. Severe underperformance in sentiment analysis. Your signals consistently lead to losses. Re-evaluate your sentiment analysis logic and filter criteria.
DECISION: Regret shows how often you defeated the returns by baseline, returns shows how much better or worse your results were, and Sharpe ratio shows how much better or worse your risk-adjusted performance was, each calculated as a percentage difference from the baseline. Regret (Outperformance): 100.0. Your decisions dominated the baseline, delivering superior returns with balanced risk. Maintain your current strategy but remain vigilant for changing conditions. Sharpe Ratio: -72.07%. Severe risk in decisions. Your allocations are erratic and consistently lead to losses. Redesign your decision logic. Returns: -71.19%. Severe underperformance in decision-making. Your allocations consistently lose value. Re-evaluate your decision logic entirely.
This fixed weekly prompt gave each agent a stable interpretation of the three core metrics: regret, returns, and Sharpe ratio. The prompt stated that regret was 100% for all agents, which indicated that they were consistently beating the baseline portfolio. The same prompt also reported strongly negative returns and Sharpe values, showing that the losses came from volatility and allocation size rather than from incorrect market direction. Keeping this message constant for the entire week meant that the agents did not redefine their objectives widely in response to changing daily feedback. This stability allowed them to update their internal decision rules gradually and consistently toward the same conclusion: maintain directional accuracy while reducing unnecessary allocation variability.
This design created an asymmetric effect across agents. The Signals agent, whose reasoning depends on interpreting natural language (news flow, sentiment reports, and textual priors), was able to incorporate such guidance effectively. For example, phrases such as “your signals consistently beat the baseline” reinforced its conviction, while instructions to prioritise high-confidence sources translated directly into more decisive sentiment weighting. This alignment between the form of feedback and the nature of the task explains why the Signals Agent produced a cumulative gain of approximately 6% during the week in which the long-term prompt remained unchanged, compared with the 4% gain observed in the same period without stable weekly guidance.
@@ 7 Contributions and Conclusion
In sum, this paper shows that an adaptive, feedback-driven multi-agent system can generate consistent Bitcoin trading profits under volatile conditions. By prioritizing profitability over predictive accuracy and employing language-based feedback as its optimization mechanism, the system demonstrates a scalable pathway for deploying LLMs in financial trading.